### v2

@@ Abstract
Since its advent in 2011, boson-sampling has been a preferred candidate for demonstrating quantum advantage because of its simplicity and near-term requirements compared to other quantum algorithms. We propose to use a variant, called coarse-grained boson-sampling (CGBS), as a quantum Proof-of-Work (PoW) scheme for blockchain consensus. The users perform boson-sampling using input states that depend on the current block information and commit their samples to the network. Afterward, CGBS strategies are determined which can be used to both validate samples and to reward successful miners. By combining rewards to miners committing honest samples together with penalties to miners committing dishonest samples, a Nash equilibrium is found that incentivizes honest nodes. The scheme works for both Fock state boson sampling and Gaussian boson sampling and provides dramatic speedup and energy savings relative to computation by classical hardware.
@@
Blockchain technology relies on the ability of a network of non-cooperating participants to reach consensus on validating and verifying a new set of block-bundled transactions, in a setting without centralized authority. A consensus algorithm is a procedure through which all the peers of the blockchain network reach a common agreement about the present state of the distributed ledger. One of the best tested consensus algorithms which have demonstrated robustness and security is Proof-of-Work (PoW) [1]. PoW relies on validating a proposed block of new transactions to be added to the blockchain by selecting and rewarding a successful “miner” who is the first to solve a computational puzzle. This puzzle involves a one-way function, i.e. a function that is easy to compute, and hence easy to verify, but hard to invert. Traditionally the chosen function is the inverse hashing problem, which by its structure makes the parameters of the problem dependent on the current block information, thus making pre-computation infeasible. Additionally, the problem is progress-free, meaning the probability of successfully mining a block at any given instant is independent of prior mining attempts. This means a miner’s success probability essentially grows linearly with the time spent, or equivalently work expended, solving the problem. The latter feature ensures that the mining advantage is proportionate to a miner’s hashing power.
@@
There are, however, two issues that threaten to compromise the continued usage of PoW consensus in a scalable manner. The first is energy consumption. Problems like inverse hashing admit fast processing, now at speeds of tens of THash/s, by application-specific integrated circuits (ASICs). Unfortunately, the tremendous speed of these devices comes at the cost of large power consumption, and as the hashing power of the network grows so grows the energy cost per transaction. The reason is for asset-based cryptocurrencies like Bitcoin, as the overall network hashing power grows, the difficulty of the one-way function is increased to maintain a constant transaction speed. Since new bitcoins are introduced through the mining process, a constant transaction speed is desirable to maintain stability and to avoid inflationary pressures. As of May 2023, a single Bitcoin transaction had the equivalent energy consumption of an average U.S. household over 19.1 days (Digiconomist).
@@
The second issue is that PoW assumes only classical computers are available as mining resources. Quantum computing technology, while only at the prototype stage now, is rapidly developing. Quantum computers running Grover’s search algorithm [3], can achieve a quadratic speedup in solving unstructured problems like inverting one-way functions. This means if they were integrated into PoW, the progress-free condition would no longer apply and the probability of solving the problem grows super-linearly with computational time spent  [^1] Specifically the probability to solve in time $t$ grows like $p(t)=\sin^{2}(ct)$, where $c=O(\sqrt{D/H})$, $H$ is the size of the search domain for the one-way function, and $D$ is the number of satisfying arguments.. This can distort network dynamics in a variety of ways. One example [4] is the case of an isolated quantum miner competing against classical miners. Normally, after a block has been successfully won but not verified by the entire network a miner who receives this block would cease working on it and start mining a new block on top of it. Instead, a quantum miner in the middle of a Grover search can choose to measure their register mid computation, and if they succeed, broadcast the answer to the network. Because of imperfect connectivity, their solution might be accepted by a majority of the nodes instead of the first solver. Another scenario is a network with all quantum miners. Because of partial progress, the miners have a chance to win early and a mixed Nash equilibrium strategy results which favours a probabilistic distribution of times to measure [5]. This can introduce substantial fluctuations in the time to verify transactions. Workarounds can be found, such as requiring the miners to commit to a time to mine before they begin [4] or using random beacons that interrupt the search progress of quantum computers by periodically announcing new puzzles to be solved [6]. However, time stamps can be spoofed and as quantum computers speed up and are parallelized, the frequency of interrupting beacons will need to increase to avoid distortions in the consensus dynamics. A future-proofed consensus algorithm should take quantum processing into account as a core resource.
@@
We propose a new PoW consensus protocol based on boson-sampling. Boson-sampling was originally developed to demonstrate quantum supremacy, owing to its reduced resource requirements compared to the other quantum algorithms [7]. Boson-samplers are specialized photonic devices that are restricted in the sense that they are neither capable of universal quantum computing nor error correctable, through proposals have been made to find practical applications in chemistry, many-body physics, and computer science [8]. We formulate a practical application of a boson-sampling variant called coarse-grained boson-sampling (CGBS) [9, 10]. This scheme involves the equal-size grouping of the output statistics of a boson-sampler into a fixed number of bins according to some announced binning tactic.
@@
Photonic-based blockchain has been investigated before. Optical PoW [11] uses HeavyHash, a slight modification of the Bitcoin protocol, where a photonic mesh-based matrix–vector product is inserted in the middle of mining. This has already been integrated into the cryptocurrencies optical Bitcoin and Kaspa. Recently, a more time and energy-efficient variant named LightHash has been tested on networks of up to $4$ photons [12]. Both of these protocols use passive linear optics networks acting upon coherent state inputs which implement matrix multiplication on the vector of coherent amplitudes. It is conjectured that the photonic implementation of this matrix multiplication can achieve an order of magnitude speedup over traditional CPU hardware. They exploit the classical speedup associated with a photonic implementation of this operation and do not exploit any quantum advantage. While that method uses a multi-mode interferometer similar to what we describe in this work, it does not use intrinsically quantum states of light and in fact, is a different form of classical computing using light. In contrast, our boson sampling method uses quantum resources with processes that become exponentially harder, in the number of photons, to simulate with classical hardware whether photonic or not.
@@ I.1.1 Fundamentals
The linear optics transformation $U$ is chosen uniformly at random from the Haar measure, essential to the underlying theoretical complexity proof. It was shown by [18] that any $M\times M$ linear optics transformation of the form shown in Eq. 2 can be decomposed into a network of at most $O(M^{2})$ beamsplitters and phase-shifters, ensuring that efficient physical implementation is always possible. As presented in Fig. 1, the number of detectors equals the number of modes $M$. In practice, the number of detectors can be reduced by exploiting multiplexing in other degrees of freedom, such as the temporal degree of freedom. For example, in the architecture presented in Ref. [19], where modes are encoded temporally, a single time-resolved detector is sufficient for detecting and distinguishing between all modes.
@@ I.1.2 Mode-binned boson-sampling
By Eq. 12, the mode-binned probability distribution can be computed by evaluating $\binom{N+d^{(\mathtt{mb})}-1}{N}$ permanents. To exactly compute the permanent of an $N\times N$ matrix requires $O(N2^{N})$ elementary operations using Ryser’s algorithm, but if we only demand a polynomial additive approximation then a cheaper computational method is available. We can use the Gurvitz approximation which allows for classical estimation of the permanent of a complex $N\times N$ matrix to within additive error $\delta$ in $O(N^{2}/\delta^{2})$ operations. The algorithm works by sampling random binary vectors and computing a Glynn estimator (Appendix A.2). The number of random samples $m$ needed to approximate $\chi({\bf s})$ to within $\delta$ with probability at least $p$ is
@@ II Results
We utilize two types of binning for PoW consensus, one for validation to catch cheaters, and one to reward miners. The former can be estimated with classical computers efficiently, while the latter does not have a known classical computation though it does have an efficient quantum estimation. Upon successful mining of a block, the output of both binning distributions will be added to the blockchain, meaning one part can be verified efficiently by classical computers while another part cannot. This will incentivize nodes using boson-sampling devices to verify prior blocks in the blockchain. The protocol is illustrated in Fig. 3 and a detailed description is provided in Sec. IV. See Table 1 in the supporting information text for a description of the various parameters. An alternate approach based on Gaussian boson-sampling is described in Sec. IV.1.
@@ II.1 Robustness
The record ${\tt Rec}(b_{j})$ can be verified since the output distribution $P^{(\mathtt{mb})}$ can be checked in polynomial time (in the number of bins $d^{(\mathtt{mb})}$ and $N$) on a classical computer. The peak probability $\mu_{\rm net}$ can be checked in polynomial time (in the number of bins $d^{(\mathtt{sb})}$) on a quantum boson-sampler. The fact that the miners don’t know the mode binning ahead of time, of which there are $M!/(M/d^{(\mathtt{mb})})!^{d^{(\mathtt{mb})}}$ possibilities, means that even after the problem is specified, there is no advantage in using even classical supercomputers to estimate $P^{(\mathtt{mb})}$. Even if the probability to randomly guess the correct mode binned distribution were non-negligible, for example, because of a choice to use a small $d^{(\mathtt{mb})}$ and large $\beta$ to speed up the validation time, provided it is smaller than $p^{\mathtt{cheat}}$, the protocol is robust. The reason is, as established in Appendix A.4, cheaters will be disincentivized since failure to pass the test incurs a penalty of lost staked tokens. Similarly, not knowing the state-binning means that they have no potential advantage in the payout.
@@
where $R_{q}$ is based on publicly available knowledge of the boson sampling repetition rate at the time of the genesis block. This choice for mining time is made to ensure that honest miners with boson samplers will have produced enough samples to pass the validation test and even if there is only one honest node, that node will have produced enough samples to earn a reward. The repetition rate will of course increase with improvements in quantum technology but that can be accommodated by varying the other parameters in the problem such as photon number, bin numbers, and prescribed accuracy, in order to maintain a stable block mining rate. For $N=25$, $d^{(\mathtt{mb})}=3$, and $\beta=0.1$, and assuming the boson sampling specs in Fig. 7, the minimum mining time would be $81.6$s. The validation test sets a lower limit on the time to execute the entire protocol and add a new block. The classical computation involved during the validation step, while tractable, can be a long computation even for moderate-sized bin numbers $d^{(\mathtt{mb})}$ and photon numbers. However, it is a problem amenable to distributed computation since the problem specification is known to all miners.
@@
In general, the protocol is immune to the most common noise sources in the photonic implementation of boson sampling. Photon loss in individual nodes’ boson sampling implementations is irrelevant since only photon number-conserving samples are committed to the network and only the sampling rate of the nodes is reduced, motivating them to improve their setups. Similarly, unital errors in individual setups only increase the variation distance between the target and erroneous distributions linearly in the number of photons and operator norm distance between the unitaries [23], which is independent of the mode numbers. In a scenario where individual nodes implement unitaries that are $\Delta$ close to the target unitary in the operator norm, the net set of committed samples appears to be a result of a stochastic unitary also $\Delta$ close to the target unitary on average, giving a good approximation of the target probability distribution. Furthermore, in the mode-binning subroutine, a suitable $\beta$ can be fixed to allow for the protocol to be run with limited partial photon distinguishability but still retaining the computational hardness of boson sampling [24]. Since sampling from partially distinguishable photons can be expressed in terms of the interference of fewer completely indistinguishable photons [25], and for large photon indistinguishability, the sampling distribution is still hard to approximate classically.

### v3

@@ Abstract
Since its advent in 2011, boson sampling has been a preferred candidate for demonstrating quantum advantage because of its simplicity and near-term requirements compared to other quantum algorithms. We propose to use a variant, called coarse-grained boson-sampling (CGBS), as a quantum Proof-of-Work (PoW) scheme for blockchain consensus. The users perform boson sampling using input states that depend on the current block information and commit their samples to the network. Afterwards, CGBS strategies are determined which can be used to both validate samples and reward successful miners. By combining rewards for miners committing honest samples together with penalties for miners committing dishonest samples, a Nash equilibrium is found that incentivizes honest nodes. We provide numerical evidence that these validation tests are hard to spoof classically without knowing the binning scheme ahead of time and show the robustness of our protocol to small partial distinguishability of photons. The scheme works for both Fock state boson sampling and Gaussian boson sampling and provides dramatic speedup and energy savings relative to computation by classical hardware.
@@
Blockchain technology relies on the ability of a network of non-cooperating participants to reach consensus on validating and verifying a new set of block-bundled transactions, in a setting without centralized authority. A consensus algorithm is a procedure through which all the peers of the blockchain network reach a common agreement about the present state of the distributed ledger. One of the best-tested consensus algorithms which have demonstrated robustness and security is Proof-of-Work (PoW) [1]. PoW relies on validating a proposed block of new transactions to be added to the blockchain by selecting and rewarding a successful “miner” who is the first to solve a computational puzzle. This puzzle involves a one-way function, i.e. a function that is easy to compute, and hence easy to verify, but hard to invert. Traditionally the chosen function is the inverse hashing problem, which by its structure makes the parameters of the problem dependent on the current block information, thus making pre-computation infeasible. Additionally, the problem is progress-free, meaning the probability of successfully mining a block at any given instant is independent of prior mining attempts. This means a miner’s success probability essentially grows linearly with the time spent, or equivalently work expended, solving the problem. The latter feature ensures that the mining advantage is proportionate to a miner’s hashing power.
@@
There are, however, two issues that threaten to compromise the continued usage of PoW consensus in a scalable manner. The first is energy consumption. Problems like inverse hashing admit fast processing, now at speeds of tens of THash/s, by application-specific integrated circuits (ASICs). Unfortunately, the tremendous speed of these devices comes at the cost of large power consumption, and as the hashing power of the network grows, so does the energy cost per transaction. The reason is for asset-based cryptocurrencies like Bitcoin, as the overall network hashing power grows, the difficulty of the one-way function is increased to maintain a constant transaction speed. Since new bitcoins are introduced through the mining process, a constant transaction speed is desirable to maintain stability and avoid inflationary pressures. As of May 2023, a single Bitcoin transaction had the equivalent energy consumption of an average U.S. household over 19.1 days (Digiconomist).
@@
The second issue is that PoW assumes only classical computers are available as mining resources. Quantum computing technology, while only at the prototype stage now, is rapidly developing. Quantum computers running Grover’s search algorithm [3], can achieve a quadratic speedup in solving unstructured problems like inverting one-way functions. This means if they were integrated into PoW, the progress-free condition would no longer apply and the probability of solving the problem grows super-linearly with computational time spent  [^1] Specifically the probability to solve in time $t$ grows like $p(t)=\sin^{2}(ct)$, where $c=O(\sqrt{D/H})$, $H$ is the size of the search domain for the one-way function, and $D$ is the number of satisfying arguments.. This can distort network dynamics in a variety of ways. One example [4] is the case of an isolated quantum miner competing against classical miners. Normally, after a block has been successfully won but not verified by the entire network a miner who receives this block would cease working on it and start mining a new block on top of it. Instead, a quantum miner in the middle of a Grover search can choose to measure their register mid-computation, and if they succeed, broadcast the answer to the network. Because of imperfect connectivity, their solution might be accepted by a majority of the nodes instead of the first solver. Another scenario is a network with all quantum miners. Because of partial progress, the miners have a chance to win early and a mixed Nash equilibrium strategy results which favours a probabilistic distribution of times to measure [5]. This can introduce substantial fluctuations in the time to verify transactions. Workarounds can be found, such as requiring the miners to commit to a time to mine before they begin [4] or using random beacons that interrupt the search progress of quantum computers by periodically announcing new puzzles to be solved [6]. However, time stamps can be spoofed and as quantum computers speed up and are parallelized, the frequency of interrupting beacons will need to increase to avoid distortions in the consensus dynamics. A future-proofed consensus algorithm should take quantum processing into account as a core resource.
@@
We propose a new PoW consensus protocol based on boson sampling. Boson-sampling was originally developed to demonstrate quantum supremacy, owing to its reduced resource requirements compared to the other quantum algorithms [7]. Boson-samplers are specialized photonic devices that are restricted in the sense that they are neither capable of universal quantum computing nor error correctable, through proposals have been made to find practical applications in chemistry, many-body physics, and computer science [8]. We formulate a practical application of a boson-sampling variant called coarse-grained boson-sampling (CGBS) [9, 10]. This scheme involves the equal-size grouping of the output statistics of a boson sampler into a fixed number of bins according to some announced binning tactic.
@@
Photonic-based blockchain has been investigated before. Optical PoW [11] uses HeavyHash, a slight modification of the Bitcoin protocol, where a photonic mesh-based matrix-vector product is inserted in the middle of mining. This has already been integrated into the cryptocurrencies optical Bitcoin and Kaspa. Recently, a more time and energy-efficient variant named LightHash has been tested on networks of up to $4$ photons [12]. Both of these protocols use passive linear optics networks acting upon coherent state inputs which implement matrix multiplication on the vector of coherent amplitudes. It is conjectured that the photonic implementation of this matrix multiplication can achieve an order of magnitude speedup over traditional CPU hardware. They exploit the classical speedup associated with a photonic implementation of this operation and do not exploit any quantum advantage. While that method uses a multi-mode interferometer similar to what we describe in this work, it does not use intrinsically quantum states of light and in fact, is a different form of classical computing using light. In contrast, our boson sampling method uses quantum resources with processes that become exponentially harder, in the number of photons, to simulate with classical hardware whether photonic or not.
@@ I.1.1 Fundamentals
The linear optics transformation $U$ is chosen uniformly at random from the Haar measure, which is essential to the underlying theoretical complexity proof. It was shown by [18] that any $M\times M$ linear optics transformation of the form shown in Eq. 2 can be decomposed into a network of at most $O(M^{2})$ beamsplitters and phase-shifters, ensuring that efficient physical implementation is always possible. As presented in Fig. 1, the number of detectors equals the number of modes $M$. In practice, the number of detectors can be reduced by exploiting multiplexing in other degrees of freedom, such as the temporal degree of freedom. For example, in the architecture presented in Ref. [19], where modes are encoded temporally, a single time-resolved detector is sufficient for detecting and distinguishing between all modes.
@@ I.1.2 Mode-binned boson-sampling
By Eq. 12, the mode-binned probability distribution can be computed by evaluating $\binom{N+d^{(\mathtt{mb})}-1}{N}$ permanents. To exactly compute the permanent of an $N\times N$ matrix requires $O(N2^{N})$ elementary operations using Ryser’s algorithm, but if we only demand a polynomial additive approximation then a cheaper computational method is available. We can use the Gurvits’ approximation which allows for classical estimation of the permanent of a complex $N\times N$ matrix to within additive error $\delta$ in $O(N^{2}/\delta^{2})$ operations. The algorithm works by sampling random binary vectors and computing a Glynn estimator (Appendix A.2). The number of random samples $m$ needed to approximate $\chi({\bf s})$ to within $\delta$ with probability at least $p$ is
@@
In practice the number of samples required is much smaller as shown in Fig. 4b. See Sec. II.1 for more detailed discussions on this.
@@ II Results
We utilize two types of binning for PoW consensus, one for validation to catch cheaters, and one to reward miners. The former can be estimated with classical computers efficiently, while the latter does not have a known classical computation though it does have an efficient quantum estimation. Upon successful mining of a block, the output of both binning distributions will be added to the blockchain, meaning one part can be verified efficiently by classical computers while another part cannot. This will incentivize nodes using boson-sampling devices to verify prior blocks in the blockchain. The protocol is illustrated in Fig. 3 and a detailed description is provided in Sec. IV. See Table 1 in the supporting information text for a description of the various parameters. An alternate approach based on Gaussian boson sampling is described in Sec. IV.1.
@@ II.1 Robustness
The record ${\tt Rec}(b_{j})$ can be verified since the output distribution $P^{(\mathtt{mb})}$ can be checked in polynomial time (in the number of bins $d^{(\mathtt{mb})}$ and $N$) on a classical computer. The peak probability $\mu_{\rm net}$ can be checked in polynomial time (in the number of bins $d^{(\mathtt{sb})}$) on a quantum boson-sampler. The fact that the miners don’t know the mode binning ahead of time, of which there are $M!/(M/d^{(\mathtt{mb})})!^{d^{(\mathtt{mb})}}$ possibilities means that even after the problem is specified, there is no advantage in using even classical supercomputers to estimate $P^{(\mathtt{mb})}$. Even if the probability to randomly guess the correct mode binned distribution were non-negligible, for example, because of a choice to use a small $d^{(\mathtt{mb})}$ and large $\beta$ to speed up the validation time, provided it is smaller than $p^{\mathtt{cheat}}$, the protocol is robust. The reason is, as established in Appendix A.4, cheaters will be disincentivized since failure to pass the test incurs a penalty of lost staked tokens. Similarly, not knowing the state-binning means that they have no potential advantage in the payout.
@@
where $R_{q}$ is based on publicly available knowledge of the boson sampling repetition rate at the time of the genesis block. This choice for mining time is made to ensure that honest miners with boson samplers will have produced enough samples to pass the validation test and even if there is only one honest node, that node will have produced enough samples to earn a reward. The repetition rate will of course increase with improvements in quantum technology but that can be accommodated by varying the other parameters in the problem such as photon number, bin numbers, and prescribed accuracy, in order to maintain a stable block mining rate. For $N=25$, $d^{(\mathtt{mb})}=3$, and $\beta=0.1$, and assuming the boson sampling specifications in the caption of Fig. 7, the minimum mining time would be $81.6$s. The validation test sets a lower limit on the time to execute the entire protocol and add a new block. The classical computation involved during the validation step, while tractable, can be a long computation even for moderate-sized bin numbers $d^{(\mathtt{mb})}$ and photon numbers. However, it is a problem amenable to distributed computation since the problem specification is known to all miners.
@@
In general, the protocol is immune to the most common noise sources in the photonic implementation of boson sampling. Photon loss would not corrupt a node’s sample set since only photon number-conserving samples would be committed to the network. The main consequence of loss in an individual node’s boson sampler would be to reduce the sampling rate of that node and motivate them to improve their setup. Similarly, unital errors in individual setups only increase the variation distance between the target and erroneous distributions linearly in the number of photons and operator norm distance between the unitaries [23], which is independent of the mode numbers. In a scenario where individual nodes implement unitaries that are $\Delta$ close to the target unitary in the operator norm, the net set of committed samples appears to be a result of a stochastic unitary also $\Delta$ close to the target unitary on average, giving a good approximation of the target probability distribution. Furthermore, in the mode-binning subroutine, a suitable $\beta$ can be fixed to allow for the protocol to be run with partial photon distinguishability but still retaining the computational hardness of boson sampling [24]. Since sampling from partially distinguishable photons can be expressed in terms of the interference of fewer completely indistinguishable photons [25], for large enough photon indistinguishability, the sampling distribution is still hard to approximate classically. Specifically, the complexity of the classical algorithm in [25] scales as $N^{2k}2^{k}k$ for the $k^{\mathrm{th}}$ order approximation of the permanents, which would not be efficient since a value of $k$ close to $N$ would be required for the level of indistinguishability assumed in Sec. II.3 and already achieved in experiments [26]. In Fig. 5, we show how the TVD varies with the indistinguishability parameter defined in Ref [25]. As an example (Fig. 4 in Ref [25] ) for $N=25$, an accuracy of $0.01$, and indistinguishability of about 0.85 will make the algorithm exponential in its time complexity. If we focus on the same level of indistinguishability in the inset plot in Fig. 5 the TVD is roughly $0.022$. A suitable $\beta$, as discussed above could be chosen such that we accept samples that are closer than this. This is also consistent with the results we present later regarding sensitivity to binning strategy and other classical spoofing methods.
In our analysis, we aim for robustness of two kinds: (a) sensitivity to binning strategy and (b) security against Haar-averaged distribution. To analyze the sensitivity of the binning strategy, we collect the samples and calculate the sample distribution for a reference binning strategy. Then we randomly permute the output modes to simulate random binning strategies with the same bin sizes. Note that this numerical simulation can be done on the same sample set as the different binnings are sortings of the mode labels of measured outcomes. We then calculate TVD between the sample distributions of 100 such random permutations and the exact mode-binned distribution with the reference binning strategy. The variance and mean of these TVDs are plotted versus sample size and are shown in Figure 4a. It is evident from the plot that the mean of TVDs is well-gapped when compared to the sample distribution with reference binning strategy. This shows that the coarse-grained (marginalized) distribution is sensitive to the binning strategy. This also provides evidence that there exists a $\beta$ such that any classically generated distribution without the knowledge of binning strategy will be at least $\beta$ away in TVD from some reference binned distribution. A naive choice of this $\beta$ would be half of the gap between the sample distribution of reference binning strategy and the mean of permuted ones in Figure 4a. This gives us enough confidence to rule out any classical spoofing attack that can well approximate the coarse-grained distribution without prior knowledge of the binning strategy.
Next, we analyze the robustness of our mode-binning validation against potential attacks where a miner submits samples using the mode-binned probability distribution averaged over the Haar-random unitary matrix representing the interferometer [27, 28, 20]. This distribution can be computed classically by cheaters before the target unitary is even broadcast to the network. These mode-binned distributions are given for the completely distinguishable ($P_{D}$) and the completely indistinguishable or bosonic case ($P_{B}$) as follows:
$P_{D}(\textbf{k})=\frac{n!}{\prod_{z=1}^{d^{(\mathtt{mb})}}k_{z}!}\prod_{z=1}^{d^{(\mathtt{mb})}}q_{z}^{k_{z}}!$ (36)
$P_{B}(\textbf{k})=P_{D}(k)\frac{\prod_{z=1}^{d^{(\mathtt{mb})}}(\prod_{l=0}^{k_{z}-1}[1+l/k_{z}])}{\prod_{l=0}^{n-1}[1+l/m]}$ (37)
where, $k_{z}$ is the bin size which we allow in principle to vary, and $q_{z}$ = $k_{z}/m$ is the relative bin size. Ideally one would like to find a lower bound on the average case, i.e. averaged over Haar-random interferometers, TVD between the reference distribution $P^{({\rm mb})}$ and the pre-computable distributions $P_{D},P_{B}$. In the absence of such a bound, we resort to numerics. Fig. 4a shows that the TVD between the reference and Haar-averaged distribution is large enough to distinguish between them with a reasonable number of samples for 16 photons and 256 modes. The inset shows a similar trend for different photon numbers.
For the purpose of these simulations, we calculated the mode-binned distribution exactly and did not use Gurvits’ approximation as in [20]. We modified the algorithm to only use $\binom{N+d^{(\mathrm{mb})}-1}{N}$ discrete (non-uniform) points on the grid for the inverse Fourier transform instead of all the $(N+1)^{d^{(\mathrm{mb})}}$ points. This provides a reduction in the run times by over three orders of magnitude [29]. For large $N$, one would need to use Gurvits’ approximation because of the exponential complexity of the exact computation of the permanent.
For completeness, we also analyse the minimum number of samples needed to ensure the samples are validated for a particular $\beta$. This means that the TVD between the sample distribution and the exact mode-binned distribution should be less than $\beta$. We find this numerically from the plot of TVD vs. sample sizes. The required number of samples is then plotted against $N$ for various values of $\beta$ and is shown in Fig. 4b. For comparison, we include the theoretical upper bound given by Eq. 25. As an example, with $N=16$ and $\beta=0.02$, Eq. 25 suggests 14.66 billion samples are needed, whereas in practice, only 3.74 million samples are required. This demonstrates that, in practical scenarios, players need to commit far fewer samples than the theoretical upper bound to pass the validation test.
Figure 4: Performance of the protocol. (a) Dependence of the total variation distance (TVD) between the exact mode-binned distributions Eq. 12 and distributions obtained from a finite number of samples for $N=16$ photons generated from the following: the Clifford sampler in BosomSampling.jl package [30] (blue line), the average distribution in the distinguishable case as given by Eq. 36 (green line), and the average distribution for the indistinguishable case as given in Eq. 37 (purple line). The red line represents the average of the TVD between the exact distribution and the mode-binned distributions generated from a hundred random permutations of the output binning scheme. The inset shows the scaling of the TVD with different photon numbers for 10 million samples. (b) The estimated number of samples for which the TVD between the sample distribution and the exact mode-binned distribution becomes less than a fixed value ($\beta$) Vs N. This is plotted for various $\beta$. The black dashed line represents the theoretical bound given by Eq. 25 for $\beta=0.02$. All plots assume a 256-mode interferometer with six bins $(M=256\text{ and }d^{(mb)}=6)$.
Figure 5: Plot of total variation distance (TVD) between the exact mode-binned distribution and finite sample distribution as a function of the number of samples for different indistinguishability values (x). The indistinguishability parameter ranges from x = 0 (fully distinguishable photons) to x = 1 (fully indistinguishable photons). The inset plot shows how TVD varies with the indistinguishability parameter for a fixed sample size of 10 million. These simulations use $N=16,M=256$, and $d^{(\text{mb})}=6$. This shows a clear separation of the target from the distribution obtained from fully distinguishable photons and also enables us to select a suitable TVD to reject imperfect samples but accept samples with sufficiently high indistinguishability ($\approx 0.85$) which has been achieved experimentally.